Reports 1-1 of 1 Clear search Modify search
Injection system (General activities)
derossi, fiori, janssens, paoletti, chiummo - 17:14 Friday 28 February 2020 (48586) Print this report
investigations about 101 and 157 Hz peaks

This morning we performed some tapping tests and noise injections on the east flange of the IB tower, where the beam dump to dump the light coming from the IPC2 on SIB1 is located (it was not installed on the suspended bench because it is water cooled).

Here a list of the performed tests:

 

type of noise parameters where GPS start GPS end observations
tapping alog Y (vertical) direction beam dump 1266915123 1266915165  
tapping along z (beam) direction beam dump 1266915312   unlock (PRCL oscillation?)
shaker - noise generator app: brown noise 90-110 Hz volume = 20% beam dump flange 1266916656 1266916762  
 - volume = 30% - 1266916762 1266916996 see x2 in BD ACC, and Hrec 101Hz peak goes up x3 about
150-170 Hz volume = 20% - 1266917056 1266917170  
- volume = 40% - 1266917170 1266917302  
- volume = 60% - 1266917302 1266917379 see 163Hz in Hrec
- volume = 80% - 1266917379 1266917561 nice: see seismic noise in SIB2_Z and little on IB_BS_ACC_Z
sine at 157 Hz     1266917778 1266917882 see line in accelerometers, but not in Hrec
  increased (about half of volume range)   1266917882 1266918006 no see little in Hrec (SNR 2) while accelerometers see SNR x5 (BD) and x20 (SIB2)
  increased to 3/4 of full range   1266918006 1266918149  
sine at 158 Hz same volume   1266918165 1266918284  
sine at 156 Hz same volume   1266918316 2 min duration see nice in Hrec: SNR of 4?
sine at 155 Hz same volume   1266918470 1266918580 little in Hrec (x2?)
brown noise 150-160 Hz max volume 100%   1266918627 1266918965 see x3 in DB acc, x10 in SIB2 acc, and something in Hrec...
brown noise 90-110 Hz max volume 100%   1266919101   we unlocked
tapping z direction flange 1266921506 1266921580  
- z direction beam dump 1266921622 1266921680  
- x direction flange 1266921846 1266921927  
- x direction BD 1266921984 1266922057  
  y flange 1266922154 1266922234  
  y BD 1266922306 1266922377  
brown 40-60 Hz 50%   1266922877 1266923057 see in BD ACC, but not in Hrec
tapping sud pipe between IB tower and SIB2 middle 1266923353 1266923415 no effect in Hrec
- nord - 1266923440 1266923499 no effect in Hrec

 

We will post more details and the analysis as soon as possible, but for the moment what we observed is that:

  • it is difficult to state where it is the origin of the peaks because when we do tapping we excite both the flange and the beam dump
  • the 101 Hz seems to be more sensitive to the noise injection (we unlocked when we put the same kind of noise than around 157 Hz)
  • when we tapped on the pipes we did't see anything, but maybe we should have tapped also on both sides of the bellows
Comments to this report:
derossi, fiori, chiummo - 16:45 Wednesday 04 March 2020 (48615) Print this report

In the first attached file is it shown where we performed the various injections. From the analysis we can assert that:

  • We couldn't pinpoint the scatterer to a specified location, but we could localize it to the east flange of the IB tower. In fact, when we did the tapping test on the beam dump and on the flange, both the accelerometer on the beam dump (BD ACC) and the one on the inner side of the flange (IB ACC X) see the same amount of noise. The 101 Hz peak came up, instead the 157 Hz peak was not excited (see plots 2 to 4, along x, y and z directions). Also while doing the injections with the shaker on the beam dump we excited both the beam dump and the flange, as shown on plot 5. To be noticed that this time also the structure at around 157 Hz has grown in Hrec.
  • During the broadband injections we also observed an increase of coherence between PRCL and PMC TRA at 101 Hz, which already had been pointed out by the BRUCO even during quiet periods (look at plot 6). This could be explained through the PSTAB loop, which tries to compensate at those frequencies (in fact this increase in the coerence is also visible on the PSTAB error signal). However, we do not see any effect during the injections in the band 150-160 Hz.
  • Plot 7 shows the injections of lines at 155, 156, 157 and 158 Hz that had been performed with the shaker installed on the beam dump. Only the 156 and 157 Hz are coupled in PRCL and Hrec.
  • Plot 8 shows the tapping performed on the pipes linking SIB1 to SIB2. There is no coupling to Hrec. It is not surprising because there are the bellows to isolate the seismic noise. We should have done it on both sides of the bellows as well in order to verify if it is coming from the viewports.

Another test we could do is to increase the power going to the beam dump by turning the waveplates of the IPC1 (EIB) and 2 (SIB1). This would require a dedicated shift.

 

Images attached to this comment
mwas - 18:50 Wednesday 04 March 2020 (48616) Print this report

Looking at the tapping figures (2-4), especially the tapping in the z direction (figure 4), there is a broad bump between 120Hz and 140Hz that is increased in h(t) and in PRCL. If further studies are done with a shaker, then a noise injection in this band should also be done. That broadband bump, if present at lower level in normal operations, might be costing much more range than the clearly visible 101Hz and 157Hz peaks themselves.

Figure 1, shows that during the tapping (purple) a mechanical resonance at ~130Hz is excited as witnessed by the accelerometer. The excited motion is about a factor 10 above the normal conditions (blue). This correspond to a noise in h(t) that is a factor 3 above normal condition, so naively one could expect this noise to be  only a factor 3 below the sensitivity curve in normal conditions.

Images attached to this comment
derossi, chiummo - 18:12 Thursday 05 March 2020 (48631) Print this report

I tried to make the projection. Since there is no coherence between the accelerometer and Hrec I computed the coupling function and set a threshold of 2 on Hrec to distinguish between measured values and the upper limit. The measured values are the yellow points in the 1st plot, so the projection is meaningful only in the region 125-140 Hz

The 2nd plot shows the projection, where we can see that we are a factor 2 at worst (at 137 Hz) below the sensitivity.

Images attached to this comment
Search Help
×

Warning

×